OpenStudy (anonymous):

a(-3.4) b(2.4) c(4.-2) d(-3.-2) I J K L half of [ab] [dc] [ad] [bc] respectively The question is in the replies

OpenStudy (anonymous):

G : (A.1) (B.1) (C.1) (D.1) give group of points that : \[\left| \right|MA+MB+MC+MD\left| \right|=\left| \right|MB-MD\left| \right|\]

OpenStudy (anonymous):

Can you give a diagram that shows where M is?

OpenStudy (anonymous):

m is group of point , that what i am looking for .......

OpenStudy (anonymous):



OpenStudy (anonymous):

Sorry, but what is a "group of point"? a point? or a group of 4 points?

OpenStudy (anonymous):

group of points

OpenStudy (anonymous):

A group of points so that the distance from these group of points follow the relationship above?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

MA+MB+MC+MD=MB-MD MA+MC+2MD=0 this is impossible. distances are always positive.